OpenStudy (anonymous):

completing the square x^2-14x+3=0

OpenStudy (callisto):

\[ x^2-14x+3=0\]\[[ x^2-14x + (\frac{14}{2})^2]-(\frac{14}{2})^2+3=0\]\[ x^2-14x +7^2-(7)^2+3=0\]\[ (x-7)^2-49+3=0\]\[ (x-7)^2-46=0\]

OpenStudy (anonymous):

awesome, thank you.. i just needed to check n make sure i had it correct.. your awesome

OpenStudy (callisto):

You're welcome :)